OpenStudy (anonymous):

how do you solve 500(1.032)^t=750?

OpenStudy (earthcitizen):

\[500(1.032)^{t}=750 \] \[516^{t}=750\] \[tlog516=\log750\] \[\therefore t =\log750/\log516\] \[t= 1.05987\]

OpenStudy (anonymous):

thanks man

OpenStudy (earthcitizen):

no p!

OpenStudy (anonymous):

i am not quite sure this is right

OpenStudy (earthcitizen):

hehe

OpenStudy (anonymous):

\[500(1.032)^t=750\] \[(1.032)^t=\frac{750}{500}=1.5\] \[t=\frac{\ln(1.5)}{\ln(1.032)}\]

OpenStudy (anonymous):

hey your right that isn't right

OpenStudy (anonymous):

i think (actually i am sure) the answer i wrote is correct

OpenStudy (anonymous):

it is i just checked the back of my book

OpenStudy (anonymous):

i get 12.87 rounded